OpenStudy (anonymous):

Simplify: (√3+√-7)(√3-√-7)

OpenStudy (primeralph):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

When we have a negative under a radical, it's necessary to pull out an i and make the stuff under the radical positive like this: \[(\sqrt{3} + \sqrt{-7})(\sqrt{3} - \sqrt{-7})\] \[(\sqrt{3} + i \sqrt{7})(\sqrt{3} - i \sqrt{7})\]From here, you can solve this by using the FOIL method, are you familiar with this method?

OpenStudy (anonymous):

yes, extremely

OpenStudy (anonymous):

Ok, so FOIL that out.

OpenStudy (primeralph):